OpenStudy (anonymous):

Please help. Will give medal! Solve the system of equations by elimination. 3x + 4y = -31 2x - 4y = -6

OpenStudy (solomonzelman):

add the equations, and tell me what you get.

OpenStudy (anonymous):

Would it be 5x = -37? I'm sorry I'm really bad at this

OpenStudy (solomonzelman):

Yes, good

OpenStudy (solomonzelman):

Now, you can solve for x

OpenStudy (anonymous):

x = 7.4?

OpenStudy (anonymous):

My answer choices are (4, 5) (5,4) (-5, 12.5) and (5, -4)

OpenStudy (solomonzelman):

x = -37/5 = -7.4

OpenStudy (anonymous):

Would I just need to plug the choices in??

OpenStudy (solomonzelman):

your shoices are wrong

OpenStudy (anonymous):

Those are the 4 answers I have to choose from

OpenStudy (solomonzelman):

They are wrong

OpenStudy (anonymous):

They can't be wrong...it's a multiple choice question

OpenStudy (solomonzelman):

We have just found that x=-7.4. This is exactly correct!! Your answers don't contain that x-coordinate, and therefore they are not correct!

OpenStudy (anonymous):

Btw I wrote the equation wrong as well...it's actually 31 positive not negative

OpenStudy (solomonzelman):

Good

OpenStudy (solomonzelman):

3x + 4y = 31 2x - 4y = -6

OpenStudy (solomonzelman):

Add the equations

OpenStudy (anonymous):

Let me try to figure it out again by myself...I'll get back to you when I do.

OpenStudy (anonymous):

Theres no way they could be "wrong". It's a multiple choice question. I have to use one of those 4

OpenStudy (anonymous):

I'm thinking I just have to plug in my answer choices and see from there

OpenStudy (solomonzelman):

For, 3x + 4y = 31 2x - 4y = -6 they are probably right. But, 3x + 4y = -31 2x - 4y = -6 as you initially wrote, they would be wrong

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

You can plug in the answer-choices, or you can solve for x and y.

OpenStudy (anonymous):

The answer was 5,4)

OpenStudy (anonymous):

3(5) + 4(4) = 31 2(5) - 4(4) = -6 15 + 16 = 31 10-16=-6 25 = 25

OpenStudy (solomonzelman):

``` The addition gives, 3x   + 4y = 31 2x - 4y = -6 ---------------- 5x - 0y = 25 5x=25 x=5 So, you can solve for y. 3x + 4y = 31 3•5 + 4y = 31 15 + 4y = 31 4y = 31 - 15 4y = 16 y = 4 So the answer is: x=5, y=4 , or, (5,4)

OpenStudy (anonymous):

Thank you! Appreciate the help